### v2

# Yanchuan Chang1    Xu Cai2    Christian S. Jensen3    Jianzhong Qi1

Abstract
Trajectory similarity is fundamental to many spatio-temporal data mining applications. Recent studies propose deep learning models to approximate conventional trajectory similarity measures, exploiting their fast inference time once trained. Although efficient inference has been reported, challenges remain in similarity approximation accuracy due to difficulties in trajectory granularity modeling and in exploiting similarity signals in the training data. To fill this gap, we propose TSMini, a highly effective trajectory similarity model with a sub-view modeling mechanism capable of learning multi-granularity trajectory patterns and a $k$ nearest neighbor-based loss that guides TSMini to learn not only absolute similarity values between trajectories but also their relative similarity ranks. Together, these two innovations enable highly accurate trajectory similarity approximation. Experiments show that TSMini can outperform the state-of-the-art models by 22% in accuracy on average when learning trajectory similarity measures. The source code of TSMini is available at https://github.com/changyanchuan/TSMini.
1 Introduction
Trajectory similarity measures quantify the similarity between two trajectories and thus play an essential role in many spatio-temporal data mining tasks and queries, such as trajectory clustering Chen et al. (2005); Wang et al. (2022), anomaly detection Laxhammar and Falkman (2013); Liu et al. (2020), and $k$-nearest neighbor ($k$NN) queries Shang et al. (2018); Wang et al. (2018a). Conventional measures Eiter and Mannila (1994); Chen et al. (2005); Ranu et al. (2015) (a.k.a. non-learned measures) are typically based on heuristic rules that match up the points in two trajectories. The similarity values are then determined by the distances between matched points. These measures often use dynamic programming to speed up the computation, which, however, remain costly for long trajectories Chang et al. (2024a).
Recent studies Yao et al. (2019); Han et al. (2021); Yang et al. (2021); Fang et al. (2022); Yao et al. (2022); Chang et al. (2023); Chen et al. (2024) adopt deep learning models to learn trajectory embeddings to accelerate similarity computation (a.k.a. learned measures). The idea is that trajectory embeddings are pre-computed offline (or computed online once) and are then used multiple times. Once embeddings are obtained, the similarity between two trajectories can be approximated by the distance between their embeddings, which can be computed efficiently.
Figure 1: Different forms of trajectory input modeling.
While high efficiency has been reported Yang et al. (2021, 2022); Chang et al. (2023), we observe two issues with the learned measures that hinder their accuracy:
(1) Difficulties in modeling the movement patterns in a trajectory at different spatial granularities: Existing solutions model an input trajectory as a sequence of points or a sequence of cells enclosing the trajectory points (see “point-based” and “cell-based” in Figure 1), or they use both input representations together. Point-based approaches consume raw GPS points that capture specific locations passed by the trajectories, while they do not reflect explicitly the relationships between the points (i.e., movement patterns). Cell-based approaches discretize the underlying space with a grid and replace GPS points by the cells enclosing them. The cell-based representation only captures rough locations passed by the trajectories, and it is difficult to determine a grid granularity that retains just the right level of detail of the trajectories. Both approaches rely on a single granularity and may not retain the key movement patterns of different trajectories.
(2) Difficulties in fully exploiting similarity signals in training data: Existing solutions minimize the mean square error (MSE) between the predicted similarity of models and the ground-truth similarity, which is typically defined by a non-learned measure. The ground-truth similarity values can come from a large continuous domain, which may not be fully reflected by the MSE over individual training samples. Thus, it is difficult for a model to learn the similarity concept in the training data from just examining the MSE results.
To address these issues, we propose TSMini, a highly effective trajectory similarity learning model with a sub-view encoder and a $k$NN-guided loss.
The sub-view encoder prepares and encodes multi-grained sub-views of input trajectories. As Figure 1 shows, a sub-view of a trajectory is a consecutive sub-sequence of the trajectory points that captures fine-grained local movement patterns. The decomposition of a trajectory into sub-views is done recursively, hence forming a series of sub-views capturing movement patterns at different granularities (Figure 3).
The output of the sub-view encoder is fed to a one-layer trajectory encoder for trajectory embedding learning. This simple encoder is a side benefit of our multi-grained sub-view modeling, as much of the pattern learning task occurs in the sub-view encoder.
To train TSMini, we employ a $k$ nearest neighbor ($k$NN)-guided loss function to better exploit training signals in the data. The intuition is that, instead of learning from individual similarity values defined by some non-learned measure (e.g., DTW Keogh and Ratanamahatana (2005)) between each trajectory pair, we examine the relative similarity between different pairs of trajectories to guide TSMini to better learn the similarity space defined by the non-learned measure. The $k$NN-guided loss achieves this by imposing a penalty when a trajectory and one of its $k$NN trajectories are predicted to be less similar than the trajectory and any of its non-$k$NNs.
To sum up, our main contributions are as follows:
(1) We propose TSMini, a highly effective model for trajectory similarity learning with a sub-view encoder and a $k$NN-guided loss.
(2) The sub-view encoder captures multi-grained movement patterns in individual trajectories, while the $k$NN-based loss guides TSMini to learn the relative similarity among multiple pairs of trajectories. Together, they enable TSMini to better learn the trajectory similarity space.
(3) We conduct extensive experiments on three large real datasets. The results show that TSMini can outperform the state-of-the-art models substantially and consistently, with an accuracy improvement of over 22% on average.
2 Related Work
Non-learned Trajectory Similarity Measures
Non-learned measures typically leverage dynamic programming or enumeration to find optimal point matches as defined by hand-crafted rules, based on which trajectory similarity values are calculated Huttenlocher et al. (1993); Eiter and Mannila (1994); Vlachos et al. (2002); Chen and Ng (2004); Chen et al. (2005); Keogh and Ratanamahatana (2005); Ranu et al. (2015). They typically have quadratic time complexity to the number of points on trajectories.
Learned Trajectory Similarity Measures
Learned measures typically first encode trajectories into embedding vectors and then compute similarity based on the embeddings.
Earlier works mainly use recurrent neural networks (RNNs) as the trajectory encoder. For example, NEUTRAJ Yao et al. (2019) encodes raw trajectory points. It introduces an MSE loss weighted by the similarity of the trajectories to focus on the most similar trajectories. TMN Yang et al. (2022) builds upon NEUTRAJ and directly takes two trajectories as input to predict a similarity value without embedding computation or reuse. KGTS Chen et al. (2024) adopts a knowledge graph to model the grid space and learns cell embeddings and it then learns trajectory embeddings from cell embeddings with a GRU Chung et al. (2014).
More recent studies use self-attention as the trajectory encoder, such as T3S Yang et al. (2021) and TrajCL Chang et al. (2023). Both models take raw GPS points and the cell sequence as the input. T3S applies a self-attention model and an LSTM Hochreiter and Schmidhuber (1997) to encode the two types of inputs, respectively, while TrajCL introduces a dual-feature self-attention module to fuse the features of raw points and grid cells. Unlike T3S and TrajCL, TSMini recursively constructs the features at multiple granularities from sub-trajectories. Another work, TrajGAT Yao et al. (2022), uses a graph-based self-attention encoder. It utilizes a quadtree Samet (1984) to partition the space into cells which form the encoder input.
Besides, TrjSR Cao et al. (2021) and ConvTraj Chang et al. (2024b) transform trajectories into 2D images and use convolutional layers to learn trajectory representation. ConvTraj also encodes original trajectory inputs. However, the generated images are mapped to specific spatial regions, which limits their generalizability, while our trajectory encoder does not suffer from such an issue. Further, these methods learn their own stand-alone measure, which differs from our setting of learning to approximate a given ground-truth measure.
Unlike these methods, we use multi-grained sub-views as the input and a $k$NN-guided loss, enabling our model to capture trajectory movement patterns at different granularity and the relative similarity between trajectories.
Figure 2: Overview of TSMini: TSMini first processes raw points on a trajectory $T$ into multi-grained sub-views and generates sub-view embeddings $\mathbf{X}$ by a Sub-view Encoder. The embeddings are fed into a Trajectory Encoder to generate embeddings $\mathbf{h}$. TSMini is optimized by the MSE loss $\mathcal{L}_{mse}$ and our $k$NN-guided loss $\mathcal{L}_{knn}$, learning both exact similarity values and relative similarity between trajectories.
3 Preliminaries
Trajectory
A trajectory $T=[p_{1},p_{2},\cdots,p_{n}]$ is a sequence of $n$ location points, where $p_{i}$ is the $i$-th point (represented by a pair of coordinates), and $\overline{p_{i}p_{j}}$ is the segment from $p_{i}$ to $p_{j}$.
Trajectory Similarity Query
Given a trajectory dataset $\mathcal{D}$, a query trajectory $T_{q}$, and query parameter $k$, and a trajectory similarity measure $f$, a trajectory similarity query (a.k.a. trajectory $k$NN query) returns a set $\mathcal{S}\subseteq\mathcal{D}$ with $|\mathcal{S}|=k$ such that $\forall T\in\mathcal{S}$ and $\forall T^{\prime}\in\mathcal{D}\setminus\mathcal{S},f(T,T_{q})\geqslant f(T^{\prime},T_{q})$.
Problem Statement
Given a trajectory similarity measure $f$ and a trajectory dataset $\mathcal{D}$, our aim is to learn a trajectory encoder model $g:T\rightarrow\textbf{h}$, where $\textbf{h}\in\mathbb{R}^{d}$ is a $d$-dimensional embedding vector, with the following two Goals:
(1) For trajectories $T_{i},T_{j}\in\mathcal{D}$, the difference between the ground-truth similarity value $f(T_{i},T_{j})$ and the predicted value by the model $f^{\prime}(T_{i},T_{j})$ is minimized, where $f^{\prime}(T_{i},T_{j})=1-\text{dist}(g(T_{i}),g(T_{j}))$ and $\text{dist}(\cdot)$ is a simple distance metric such as $L_{1}$ or $L_{2}$ norm distances.
(2) For any $T_{i}\in\mathcal{D}$ as the query trajectory, the difference between the ground-truth trajectory similarity query result set $\mathcal{S}$ and the computed result set $\mathcal{S}^{\prime}$ is minimized, where $\mathcal{S}^{\prime}$ is obtained by applying $f^{\prime}$ as the similarity measure.
Discussion
Existing studies mainly focus on Goal 1, i.e., minimizing the errors in similarity approximation. However, such a training goal does not provide sufficient signals for trajectory similarity learning. Based on the similarity values, we further exploit Goal 2 to learn the relative similarity between trajectories for more accurate similarity approximation.
4 The TSMini Model
Figure 2 illustrates our overall model structure. TSMini first encodes trajectories into sub-view embeddings through a sub-view encoder (Section 4.1) and then maps the sub-view embeddings into trajectory embeddings through a trajectory encoder (Section 4.2). Finally, TSMini leverages a $k$NN-guided loss for model training (Section 4.3).
4.1 Sub-view Encoder
Given a raw trajectory input $T$, the sub-view encoder (denoted as SVEnc) first transforms $T$ into sub-sequences, which we call sub-views, instead of modeling individual points as in previous works Yao et al. (2019); Yang et al. (2021); Chang et al. (2023). Trajectory sub-views can capture not only the spatial features of individual points but also local movement patterns. They provide multi-grained spatial features while reducing the input length for the subsequent trajectory encoder, hence leading to embeddings that better preserve trajectory features.
SVEnc encodes the sub-views of $T$ (augmented as below) into embeddings $\mathbf{X}\in\mathbb{R}^{m\times d}$, where $m$ denotes the number of sub-views and $d$ is the embedding dimensionality as before.
Figure 3: Multi-grained trajectory sub-view modeling.
Feature Preparation
Given an input raw trajectory $T$, we augment it by adding spatial features to each point. A point $p_{i}\in T$ is expanded to a vector of seven elements, including the longitude, the latitude, the lengths of segments $\overline{p_{i-1}p_{i}}$ and $\overline{p_{i}p_{i+1}}$, the angle between the $x$-axis and $\overline{p_{i-1}p_{i}}$, the angle between the $x$-axis and $\overline{p_{i}p_{i+1}}$, and the interior angle between $\overline{p_{i-1}p_{i}}$ and $\overline{p_{i}p_{i+1}}$. We use $\mathbf{S}\in\mathbb{R}^{n\times 7}$ to denote the augmented representation of $T$.
SVEnc
Figure 3 shows the structure of SVEnc. Inspired by time series representation learning Nie et al. (2023); Gruver et al. (2023), we stack three convolutional sub-modules sandwiched between two linear layers. Each convolutional sub-module contains a 1D convolutional (Conv1D) layer, a batch normalization Ioffe and Szegedy (2015) layer, and a LeakyReLU activation function. For all Conv1Ds, the kernel size $w_{k}$ is set as 3 and the stride $w_{s}$ is set as 1. The dimensionality of each layer is labeled in Figure 3.
Note that we do not explicitly create sub-views. Instead, we feed the entire trajectory $\mathbf{S}$ into SVEnc, where the sub-views are created and encoded simultaneously. The convolution kernels move along the dimension of trajectory points, which can be seen as first forming a sequence of sub-views of every $w_{k}$ points and then producing embeddings for the sub-views. The first Conv1D layer processes sub-views of points, while the second and third Conv1D layers aggregate every $w_{k}$ sub-views from the previous layer, summarizing the movement patterns of longer sub-sequences.
The output $\mathbf{X}\in\mathbb{R}^{m\times d}$ ($m<n$) of SVEnc is fed into the trajectory encoder to generate the trajectory embedding $\mathbf{h}$.
4.2 Trajectory Encoder
The trajectory encoder (denoted as TrajEnc) can be any existing sequence embedding model. We adapt the self-attention-based backbone model from Llama-2 Touvron et al. (2023) as TrajEnc for TSMini. This choice comes with two advantages: (1) Self-attention is both effective and efficient for trajectory similarity learning Yang et al. (2021); Chang et al. (2024a). (2) The self-attention-based backbone model in Llama-2 is shown to be more effective than the vanilla self-attention Vaswani et al. (2017) for capturing sequential correlation (see Appendix C.4 in supplementary material; same below).
Next, we detail TrajEnc. The right part of Figure 2 illustrates TrajEnc. The sub-view embeddings $\mathbf{X}$ first go through a root mean square normalization layer (RMSNorm) Zhang and Sennrich (2019) to obtain normalized input, for more stable model training. We abuse the notation slightly and also use $\mathbf{X}$ to denote the normalized input for conciseness, when there is no ambiguity.
Then, the normalized $\mathbf{X}$ is fed into a multi-head self-attention layer (MHSA) to learn the hidden correlations between trajectory sub-views and generate new embeddings $\mathbf{H}$:
$\begin{split}\mathbf{Q}^{j},\mathbf{K}^{j},\mathbf{V}^{j}&=\mathbf{X}\mathbf{W}_{Q}^{j},\mathbf{X}\mathbf{W}_{K}^{j},\mathbf{X}\mathbf{W}_{V}^{j},\\ \mathbf{H}^{j}&=\mathrm{Softmax}(\alpha_{A}\cdot\mathrm{RoPE}(\mathbf{Q}^{j},\mathbf{K}^{j}))\mathbf{V}^{j},\\ \mathbf{H}&=\mathrm{Concat}(\mathbf{H}^{1},\cdots,\mathbf{H}^{j},\cdots,\mathbf{H}^{h})\mathbf{W}_{H}.\end{split}$ (1)
Here, $\mathbf{W}_{Q}^{j}$, $\mathbf{W}_{K}^{j}$, and $\mathbf{W}_{V}^{j}$ (all in $\mathbb{R}^{d\times(d/h)}$) are learnable weights of the $j$-th self-attention head ($h$ heads in total), $\alpha_{A}$ is a scaling factor, and $\mathbf{W}_{H}\in\mathbb{R}^{d\times d}$. $\mathrm{RoPE}(\cdot)$ computes the attention coefficients and integrates with positional encodings. We adopt the Rotary Position Embedding (RoPE) Su et al. (2024) for its capability in exploiting the relative position dependency among the sub-views.
Then, we apply a residual connection He et al. (2016) to the output of the self-attention layer and the initial input $\mathbf{X}$ (before normalization), i.e., adding the two values. This alleviates gradient vanishing and yields smooth gradients.
Next, the sum from the residual connection is fed into another sub-module that consists of an RMSNorm, a feed forward network (FFN) with an SwiGLU activation function Shazeer (2020), and a residual connection, which brings the nonlinear representation learning capability. The output of this sub-module, $\widetilde{\mathbf{H}}$, has the same shape as $\mathbf{H}$ and $\mathbf{X}$.
All layers above are stacked (i.e., repeated) in Llama-2, while we found that such repetition is not necessary for our task (see Appendix C.5) and thus we do not stack the layers.
Finally, we apply average pooling to the hidden output $\widetilde{\mathbf{H}}$ to obtain the final trajectory embedding of $T$, i.e., $\mathbf{h}\in\mathbb{R}^{d}$.
Discussion
TrajEnc in TSMini is not exactly the self-attention model in Llama-2. TrajEnc is an encoder, while the self-attention model in Llama-2 is closer to a Transformer decoder. Computing a trajectory embedding by our model can be viewed as encoding a prompt in LLMs without the inference step. To adapt a “decoder” into an “encoder”, we made two main changes as follows:
(1) TrajEnc does not adopt the grouped-query attention (GQA) Ainslie et al. (2023) used by Llama-2 for fast inference, which shares attention coefficients among attention heads. TrajEnnc has much few parameters already and hence we do not need GQA which sacrifices accuracy for efficiency.
(2) TrajEnc does not use causal masking that prevents future token inputs from impacting attention computation in LLMs. This is because an entire trajectory is known for encoding in our task. To encode a point, it is more beneficial to consider both proceeding and subsequent (i.e., future) points on the trajectory, instead of just the preceding ones.
4.3 $K$NN-Guided Optimization
Next, we present our $k$NN-guided loss. To simplify the symbols, we abuse the notation slightly and use the same symbol to represent both a random variable and an observed value of the variable, e.g., $\mathbf{x}$, when the context is clear.
Figure 4: Variable dependency for the $k$NN-guided loss.
$K$NN-guided Loss
We formulate the $k$NN-guided loss $\mathcal{L}_{knn}$ in a probabilistic manner, as shown in Figure 4. Let $\mathbf{x}$ be a given variable for a list of size $N$ of the predicted similarity scores between a trajectory and all trajectories in the training set. Let $k\in\mathcal{K}$ be the query parameter $k$ of $k$NNs, where $\mathcal{K}=\mathbb{N}_{\leq N}$. We introduce $\phi\in\Phi$ as a hidden variable determined by $\mathbf{x}$ and $k$. Here, $\phi$ denotes a partially sorted list of size $N$, where the first $k$ elements are larger than the rest, i.e., the first $k$ elements in $\phi$ form the $k$NN result set. We use $\mathbf{y}$ to represent the ground-truth list of similarity scores of size $N$, which are impacted by the input variables $\mathbf{x}$, $k$, and $\phi$. The likelihood function for $\mathbf{y}$ conditioned on $\mathbf{x}$ is:
$P(\mathbf{y}|\mathbf{x})=\sum_{k\in\mathcal{K}}P(\mathbf{y}|\mathbf{x},k)P(k).$ (2)
Given a set of $N$ training instances $\{(x,y)\}$, where $x$ is a list of predicted similarity scores between a trajectory and all trajectories in the training trajectory set, and $y$ is the corresponding ground-truth list, the $k$NN-guided loss $\mathcal{L}_{knn}:(\mathbf{y},\mathbf{x})\rightarrow\mathbb{R}$ is designed to minimize the negative log-likelihood, as shown in Eq. (2):
$\mathcal{L}_{knn}(\mathbf{y},\mathbf{x})=-\log_{2}\sum_{k\in\mathcal{K}}P(\mathbf{y}|\mathbf{x},k)P(k).$ (3)
This log-sum loss is difficult to optimize directly due to its poor numerical stability (when the summation of probabilities stays close to 1, the loss will remain near 0 causing the gradient vanishing problem). To address this issue, recognizing the convexity of the negative logarithm function, we leverage Jensen’s inequality to derive an upper bound for $\mathcal{L}_{knn}$. Specifically, for a convex function $\psi$, Jensen’s inequality states that $\psi(\sum x_{i})\leq\sum(\psi(x_{i}))$, which allows us to construct a surrogate upper bound for Eq. (3), as shown in Eq. (4). Further, since $P(k)$ represents the probability of $k$, its value always lies between 0 and 1. Thus, the upper bound simplifies to Eq. (5), which we denote as $\mathcal{L}_{knn}^{+}$.
$\mathcal{L}_{knn}(\mathbf{y},\mathbf{x})$ $=-\log_{2}\sum_{k\in\mathcal{K}}P(\mathbf{y}|\mathbf{x},k)P(k)$
$\leqslant-\sum_{k\in\mathcal{K}}\log_{2}P(\mathbf{y}|\mathbf{x},k)P(k)$ (4)
$\leqslant-\sum_{k\in\mathcal{K}}\log_{2}P(\mathbf{y}|\mathbf{x},k).$ (5)
To proceed with Eq. (5), we expand $P(\mathbf{y}|\mathbf{x},k)$ as:
$P(\mathbf{y}|\mathbf{x},k)$ $=\frac{P(\mathbf{y},\mathbf{x},k)}{P(\mathbf{x})P(k)}=\frac{1}{P(\mathbf{x})P(k)}\sum_{\phi\in\Phi}P(\mathbf{y},\mathbf{x},k,\phi)$
$=\frac{1}{P(\mathbf{x})P(k)}\sum_{\phi\in\Phi}P(\mathbf{x})P(k)P(\phi|\mathbf{x},k)P(\mathbf{y}|\mathbf{x},k,\phi)$
$=\sum_{\phi\in\Phi}P(\phi|\mathbf{x},k)P(\mathbf{y}|\mathbf{x},k,\phi)$
$=P(\bar{\phi}|\mathbf{x},k)P(\mathbf{y}|\mathbf{x},k,\bar{\phi})+\sum_{\phi\in\Phi\setminus\bar{\phi}}P(\phi|\mathbf{x},k)P(\mathbf{y}|\mathbf{x},k,\phi)$
$=P(\mathbf{y}|\mathbf{x},k,\bar{\phi}).$ (6)
Here, in the first step, we exploit Bayes’ rule and the fact that $\mathbf{x}$ and $k$ are independent. The subsequent steps use the dependency between $\mathbf{y}$, $\mathbf{x}$, $k$, and $\phi$ as discussed at the start of this section. For the final step, we reduce the summation over all possible $\phi$ to a special $\bar{\phi}$, which represents the case where all the elements in $\phi$ are sorted in descending order. Therefore, $P(\bar{\phi}|\mathbf{x},k)=1$ and $P(\Phi\setminus\bar{\phi}|\mathbf{x},k)=0$, since we only examine sorted $\phi$ in TSMini.
By substituting Eq. (6) into Eq. (3), $\mathcal{L}_{knn}^{+}$ becomes:
$\mathcal{L}_{knn}^{+}(\mathbf{y},\mathbf{x})=-\sum_{k\in\mathcal{K}}\log_{2}P(\mathbf{y}|\mathbf{x},k,\bar{\phi}).$ (7)
We now formulate the likelihood $P(\mathbf{y}|\mathbf{x},k,\phi)$ using the Bradley-Terry model with a reward function $w(\mathbf{y})$ Bradley and Terry (1952). The Bradley-Terry model estimates the outcome of pairwise comparisons, determining the probability that one item is preferred over another. Formally,
$P(y_{i}>y_{j}|x_{i},x_{j},k,\phi)=\sigma(x_{i}-x_{j}),$ (8)
where $\sigma$ is the sigmoid function, and $x_{i}$ (or $y_{i}$) denotes the $i$-th element in $x$ (or $y$), $i\in[1,k]$ and $j\in(k,N]$. The Bradley-Terry model aligns well with the $k$NN query-based objective. They both aim to find a list where the probability that the $i$-th trajectory is more similar to the query trajectory than the $j$-th trajectory based on the predicted similarity scores.
Substituting the Bradley-Terry reward function in Eq. (8) into Eq. (7), we can further derive an upper bound for $\mathcal{L}_{knn}^{+}$.
$\mathcal{L}_{knn}^{+}(\mathbf{y},\mathbf{x})$ $=-\sum_{k\in\mathcal{K}}\log_{2}\prod_{i=1}^{k}\prod_{j=k}^{N}P(y_{i}>y_{j}|x_{i},x_{j},k,\bar{\phi})^{w(\mathbf{y})}$
$=-\sum_{k\in\mathcal{K}}\sum_{i=1}^{k}\sum_{j=k}^{N}\log_{2}P(y_{i}>y_{j}|x_{i},x_{j},k,\bar{\phi})^{w(\mathbf{y})}$
$=-N\sum_{i=1}^{k}\sum_{j=k}^{N}\log_{2}P(y_{i}>y_{j}|x_{i},x_{j},\bar{\phi})^{w(\mathbf{y})}$ (9)
$\leqslant-N\sum_{y_{i}>y_{j}}\log_{2}P(y_{i}>y_{j}|x_{i},x_{j},\bar{\phi})^{w(\mathbf{y})}.$ (10)
The term in Eq.(10) turns out to be a scaled version of the LambdaLoss Wang et al. (2018b), a loss often used in learning-to-rank problems. This connection, though not immediately obvious, implies a potential synergy between $k$NN guidance and learning-to-rank techniques.
We omitted a few steps in the derivation of Eqs. (9) and (10) due to space limit. These can be found in Appendix A.
The upper bound introduced in Eq. (10), denoted as $\mathcal{L}_{knn}^{++}$, is derived based on the constraint that for any given $k$, we compare elements only between those ranked before and after the $k$-th position. As a result, the total number of comparisons is limited to a maximum of $N^{2}$. Note also that $\mathcal{L}_{knn}^{++}$ is simpler to implement, while it still captures the essential characteristics of the $k$NN guidance.
The Overall Loss Function
We implement the reward function $w(\mathbf{y})$ in $\mathcal{L}_{knn}^{++}$ using the approximated Normalized Discounted Cumulative Gain (NDCG) (Eq. (16) by Wang et al. 2018b), which is an effective ranking metric. Combining Eq. (8) with the NDCG reward function, and substituting them into Eq. (10), we obtain the $\mathcal{L}_{knn}^{++}$ over all training samples by calculating $\mathbb{E}_{\{(x,y)\}}\mathcal{L}_{knn}^{++}(y,x)$ as follows:
$\mathcal{L}_{knn}^{++}=\mathop{\mathbb{E}}_{\{(x,y)\}}-N\sum_{y_{i}>y_{j}}\delta_{i,j}(G_{i}-G_{j})\log_{2}\sigma(x_{i}-x_{j}).$ (11)
Here, $G_{i}=\frac{2^{y_{i}}-1}{\mathrm{maxDCG}}$ with $\mathrm{maxDCG}=\sum_{i}\frac{2^{y_{i}}-1}{\log_{2}(i+1)}$, and the difference of discounts $\delta_{i,j}=\frac{1}{\log_{2}(j-i+1)}-\frac{1}{\log_{2}(j-i+2)}$.
Our overall loss function $\mathcal{L}$ combines the $k$NN-guided loss with a weighted version of the MSE loss, denoted as $\mathcal{L}_{mse}$:
$\mathcal{L}_{mse}=\displaystyle\mathop{\mathbb{E}}_{T_{i},T_{j}\in\mathcal{D}}w_{i,j}\cdot\big{(}f(T_{i},T_{j})-f^{\prime}(T_{i},T_{j})\big{)}^{2}.$ (12)
We use $f(T_{i},T_{j})$ as the weights $w_{i,j}$. The MSE loss allows TSMini to also learn from the absolute similar scores. We balance the two loss terms with parameter $\lambda\in(0,1)$:
$\mathcal{L}=\lambda\mathcal{L}_{mse}+(1-\lambda)\mathcal{L}_{knn}^{++}.$ (13)
5 Experiments
5.1 Experimental Settings
Datasets
We use three widely used large trajectory datasets: Porto Porto (2015), Xian DiDi (2018), and Germany OpenStreetMap (2013). Following previous studies Yao et al. (2019); Yang et al. (2021); Yao et al. (2022); Chang et al. (2023), we remove consecutive, duplicated points and filter out trajectories with fewer than 20 points or more than 200 points. We randomly sample 10,000 trajectories from each dataset, which is further split by 7:1:2 for training, validation, and test. More details about the datasets can be found in Appendix B.
Competitors.
We compare our TSMini model with six latest models: NeuTraj Yao et al. (2019), T3S Yang et al. (2021), TMN Yang et al. (2022), TrajGAT Yao et al. (2022), TrajCL Chang et al. (2023), and KGTS Chen et al. (2024), which have been described in Section 2.
Implementation Details
We set the embedding dimensionality $d$ to 128. Parameter $\lambda$ in the loss function (Eq. (13)) is set to 0.2. The trajectory encoder uses a one-layer self-attention. We optimize TSMini by Adam Kingma and Ba (2015) with a maximum of 40 epochs and a batch size of 128. The early stop patience is set to 10. The learning rate is initialized to 0.002 and decayed by 50% after every 15 epochs. We report the average results of 3 runs for each experiment.
For the baseline models, we use their released code and default settings, except for T3S which is not publicly available. We implement T3S following its proposal. The side length of grid cells is set as 100 meters for all baseline models.
Evaluation Metrics
Following previous studies Yao et al. (2019); Yang et al. (2021); Yao et al. (2022), we use hit ratios HR@10 and HR@50 and recall R10@50 to evaluate model performance. HR@$x$ is the overlapping ratio between the ground-truth top-$x$ results and the predicted top-$x$ trajectories in test sets. R10@50 is the recall of the ground-truth top-10 results in the predicted top-50 trajectories.
DTW EDwP Fréchet
Dataset Method HR@10 HR@50 R10@50 HR@10 HR@50 R10@50 HR@10 HR@50 R10@50 Average rank
NeuTraj 0.445 0.568 0.824 0.484 0.601 0.866 0.539 0.677 0.930 3.0
T3S 0.255 0.406 0.564 0.423 0.557 0.795 0.492 0.659 0.892 4.8
TMN 0.228 0.290 0.520 0.216 0.285 0.503 0.215 0.280 0.542 7.0
TrajGAT 0.279 0.366 0.686 0.226 0.345 0.601 0.384 0.430 0.793 5.4
TrajCL 0.415 0.572 0.791 0.546 0.668 0.910 0.584 0.715 0.938 2.3
KGTS 0.464 0.497 0.784 0.535 0.501 0.812 0.343 0.331 0.618 4.4
TSMini (ours) 0.765 0.869 0.996 0.796 0.881 0.998 0.754 0.838 0.996 1.0
Porto $\Delta$ +65% +52% +21% +46% +32% +10% +29% +17% +6%
NeuTraj 0.547 0.663 0.882 0.392 0.504 0.759 0.591 0.731 0.955 3.8
T3S 0.298 0.474 0.619 0.334 0.495 0.705 0.600 0.735 0.925 5.4
TMN 0.385 0.414 0.686 0.396 0.417 0.705 0.389 0.422 0.729 6.3
TrajGAT 0.479 0.616 0.917 0.378 0.577 0.842 0.522 0.633 0.959 3.7
TrajCL 0.444 0.625 0.815 0.427 0.556 0.799 0.634 0.732 0.948 3.7
KGTS 0.553 0.594 0.868 0.606 0.574 0.871 0.458 0.445 0.754 4.0
TSMini (ours) 0.754 0.883 0.995 0.797 0.889 0.996 0.812 0.890 0.997 1.0
Xian $\Delta$ +36% +33% +8% +31% +54% +14% +28% +21% +4%
NeuTraj 0.548 0.605 0.867 0.557 0.608 0.855 0.566 0.657 0.892 3.9
T3S 0.444 0.666 0.854 0.463 0.728 0.888 0.536 0.765 0.944 3.8
TMN 0.313 0.432 0.589 0.332 0.527 0.661 0.471 0.569 0.694 6.2
TrajGAT 0.338 0.503 0.632 0.185 0.310 0.435 0.572 0.755 0.913 5.4
TrajCL 0.684 0.761 0.993 0.678 0.809 0.980 0.691 0.856 0.990 2.0
KGTS 0.454 0.497 0.728 0.465 0.496 0.729 0.381 0.362 0.598 5.7
TSMini (ours) 0.717 0.778 0.999 0.822 0.933 0.997 0.869 0.932 0.998 1.0
Germany $\Delta$ +5% +2% +1% +21% +15% +2% +26% +9% +1%
Table 1: Overall Performance Results. Best results are in boldface, while the second best is underlined. $\Delta$ denotes the improvement of TSMini over the best baseline. The average rank is the mean of the ranks of each model across all result columns in the table.
5.2 Results
Overall Results
For each set of experiments, we follow the evaluation steps used by the competitors: (1) train the models to approximate a non-learned measure (DTW, EDwP, or Fréchet); (2) compute trajectories embeddings for a test set; (3) compute the similarity between any two trajectories in the test set based on the embeddings; and (4) compute evaluation metrics by comparing the predicted similarity values with the ground-truth (computed with the non-learned measure).
The overall results are shown in Table 1. We see that TSMini is consistently the most accurate, achieving 32%, 26%, and 7% improvements on average over the best baseline models in HR@10, HR@50, and R10@50, respectively. We made the following observations:
(1) TSMini achieves high R10@50 ($>0.99$), ensuring that it can effectively filter out irrelevant results for $k$NN queries.
(2) The second best model is TrajCL on Porto and Germany, while this varies on Xian. This is because the density of trajectories in Xian is higher than that in Porto, making learning more difficult. DTW is better approximated by NeuTraj, EDwP by KGTS, and Fréchet by TrajCL.
(3) KGTS mostly has lower (i.e., worse) HR@50 than HR@10. This is because KGTS is trained without using any non-learned measure, and such unlabeled data is hard to contribute to approximate specific non-learned measures.
(a) On Xian
(b) On Germany
Figure 5: Varying the number of points on trajectories to approximate Fréchet (OOM: out-of-memory error).
[…]
Dataset Model variant HR@10 HR@50 R10@50 HR@10 HR@50 R10@50 HR@10 HR@50 R10@50
TSMini 0.765 0.869 0.996 0.796 0.881 0.998 0.754 0.838 0.996
TSMini-w/o-S 0.589 0.725 0.936 0.540 0.680 0.901 0.381 0.554 0.776
TSMini-w/o-K 0.436 0.586 0.802 0.484 0.617 0.865 0.656 0.771 0.983
Porto TSMini-w/o-SK 0.131 0.262 0.382 0.118 0.222 0.332 0.282 0.455 0.654
TSMini 0.754 0.883 0.995 0.797 0.889 0.996 0.812 0.890 0.997
TSMini-w/o-S 0.635 0.778 0.954 0.600 0.741 0.931 0.512 0.669 0.860
TSMini-w/o-K 0.339 0.532 0.670 0.344 0.570 0.814 0.765 0.835 0.993
Xian TSMini-w/o-SK 0.255 0.468 0.611 0.120 0.270 0.381 0.390 0.570 0.771
TSMini 0.717 0.778 0.999 0.822 0.933 0.997 0.869 0.932 0.998
TSMini-w/o-S 0.578 0.746 0.938 0.585 0.813 0.963 0.700 0.855 0.996
TSMini-w/o-K 0.317 0.531 0.665 0.344 0.551 0.701 0.388 0.583 0.724
Germany TSMini-w/o-SK 0.133 0.201 0.282 0.129 0.213 0.284 0.167 0.312 0.541
[…]
Table 2: Ablation Study Results.
Varying the Number of Points on Trajectories ($n$)
We investigate the effectiveness of TSMini on longer trajectories. We randomly sample another two datasets of 10,000 trajectories from Xian and Germany (Porto does not have as many long trajectories), with 200 to 400, and 400 to 800 points on trajectories, respectively. We report HR@10 on approximating Fréchet. Similar result patterns are observed on other measures and in other metrics, and thus omitted.
Figure 5 shows that, overall, the accuracy of all methods decreases as the length of trajectories increases, since longer trajectories bring challenges in representation learning.
TSMini consistently outperforms all baselines, especially achieving higher HR@10 when $n=(400,800]$ compared to other methods at $n=[20,200]$. This highlights the superior capability of TSMini on learning long trajectories. TMN and TrajGAT incur out-of-memory (OOM) errors when $n=(400,800]$. This is because TMN requires computing the correlation between all point pairs in two trajectories, while TrajGAT constructs large graphs to represent long trajectories, both of which are memory-intensive.
We also study the impact of training set size, showing that TSMini is robust against it. See Appendix C.1 for the results.
Ablation Study
We compare TSMini with three model variants: (1) TSMini-w/o-S replaces the sub-view encoder SVEnc with the cell encoder used in T3S and TrajCL; (2) TSMini-w/o-K removes the $k$NN-guided loss from Eq. (13); (3) TSMini-w/o-SK uses the cell encoder and the MSE loss. We repeat the experiments as above.
As Table 2 shows, both the $k$NN-guided loss and SVEnc are important to the overall model performance. They improve the hit ratios by 88% (TSMini-w/o-K vs. TSMini) and 41% (TSMini-w/o-S vs. TSMini) in HR@10 on average, respectively. In most cases, the $k$NN-guided loss plays a more critical role, especially in learning DTW and EDwP, whereas the sub-view modeling is more effective in learning Fréchet.
We further investigate the impact of SVEnc and the self-attention module in TrajEnc. As the results in Appendix C.4 show, SVEnc is highly effective, and our adapted self-attention module outperforms the vanilla self-attention.
Parameter Study
We study the impact of batch size $N$, the hyper-parameter $\lambda$ in the loss function, and the number of self-attention layers, by repeating the experiments as before and reporting HR@10 on Xian. Full results of parameter study can be found in Appendix C.5.
(a) Batch size $N$ in Eq. (11
(b) $\lambda$ in Eq. (13
Figure 6: Parameter study results
Figure 6a shows the results on varying $N$. The HR@10 results generally increase with $N$. A larger batch size (i.e., a larger $N$ in Eq. (11)) allows TSMini to learn more relative similarity values from the $k$NN-guided loss. We use $N=128$ by default to keep inline with the baseline models.
Figure 6b shows the results on varying $\lambda$ in Eq. (13). Except when $\lambda$ is 0 or 1, $\lambda$ has a marginal impact on TSMini. Such results indicate that both loss terms contribute to the accuracy of TSMini, while using the $k$NN-guided loss alone ($\lambda=0$) is more effective than using the MSE loss ($\lambda=1$).
Model Efficiency
We further study the time and space efficiency of TSMini. Table 3 shows the results for learning to approximate Fréchet, where the reported time is end-to-end. TSMini is efficient in both space and time: (1) It is one of the models with the fewest parameters, trailing behind NeuTraj and TMN that solely rely on RNNs, and T3S that is based on self-attention without FFNs. These models have low accuracy as reported above. (2) It is the fastest in training, benefiting from the $k$NN-guided loss that leads to fast convergence. (3) It is among the fastest in inference, due to the highly efficient parallelization of self-attention computation on GPUs. TMN cannot be used as an encoder on its own (Section 2) and hence is omitted here.
[…]
TrT InfT TrT InfT TrT InfT
NeuTraj 0.12M 5,605s 2.47s 5,661s 2.96s 6,897s 2.62s
T3S 0.20M 2,246s 0.26s 2,885s 0.30s 2,373s 0.31s
TMN 0.18M 3,968s - 6,176s - 4,204s -
TrajGAT 2.84M 15,529s 2.48s 34,866s 3.64s 9,770s 3.52s
TrajCL 1.22M 2,810s 0.20s 3,006s 0.25s 2,906s 0.24s
KGTS 2.99M 2,534s 0.14s 5,600s 0.17s 4,902s 0.17s
TSMini 0.28M 1,917s 0.28s 2,605s 0.34s 1,724s 0.32s
[…]
Table 3: Space and Time Efficiency (TrT and InfT denote training time and inference time, respectively; more results in Appendix C.6)
6 Conclusion
We proposed TSMini, a simple yet highly effective trajectory similarity learning model using a sub-view encoder for input trajectory modeling and a $k$NN-guided loss for model training. TSMini can capture trajectory movement patterns at different granularity and is optimized with the similarity between a trajectory and its $k$NNs. It thus produces trajectory representations robust to a variety of similarity measures. Experimental results on large trajectory datasets show that TSMini achieves significant improvements in accuracy, outperforming the state-of-the-art models by 22% on average.
Ethical Statement
We present a model for trajectory representation learning. The model is versatile and could support different trajectory analytic tasks such as similarity queries and clustering. However, it is not meant to be used for uncovering human trajectories which might raise privacy concerns.
Appendix A Additional Steps in Eqs. (9) and (10)
We elucidate the omitted steps in the derivation of Eqs. (9) and (10). Here, we adopt a uniform assumption on $k\in[1,N]$, i.e., $P(k)=\frac{1}{N}$. Recall that in Eq. (6), we have shown the interchangeability between $P(\mathbf{y}|\mathbf{x},k,\bar{\phi})$ and $P(\mathbf{y}|\mathbf{x},k)$. Thus, we can derive:
$P(\mathbf{y}|\mathbf{x},k,\bar{\phi})$ $=P(\mathbf{y}|\mathbf{x},k)=\frac{P(\mathbf{y},\mathbf{x},k)}{P(\mathbf{x})P(k)}$
$=\frac{N}{P(\mathbf{x})}\cdot P(\mathbf{y},\mathbf{x},k)$ (14)
$=\frac{N}{P(\mathbf{x})}\cdot\frac{P(\mathbf{y},\mathbf{x})}{N}$ (15)
$=P(\mathbf{y}|\mathbf{x})=P(\mathbf{y}|\mathbf{x},\bar{\phi}),$ (16)
where Eq. (14) uses $P(k)=\frac{1}{N}$, and Eq. (15) follows from the law of total probability $P(\mathbf{y},\mathbf{x})=\sum_{k\in\mathcal{K}}P(\mathbf{y},\mathbf{x},k)=N\cdot P(\mathbf{y},\mathbf{x},k)$, given that the probabilities conditioned on $k$ are uniform with respect to any possible combination of $\mathbf{x}$ and $\mathbf{y}$. Finally, we derive Eq. (9) by aggregating the contributions from all $k$.
Appendix B Additional Details on Experimental Settings
We use three widely used large trajectory datasets:
(1) Porto Porto [2015] contains 1.7 million taxi trajectories collected from Porto, Portugal, between July 2013 and June 2014.
(2) Xian DiDi [2018] contains 2.1 million ride-hailing trajectories from Xi’an, China, in the first two weeks of October 2018.
(3) Germany OpenStreetMap [2013] contains 0.4 million trajectories publicly owned by OpenStreetMap collected across Germany before 2013.
Table 4 summarizes the dataset statistics.
[…]
#trajectories 1,372,725 900,562 420,074
Avg. #points per traj. 48 118 72
Max. #points per traj. 200 200 200
Avg. traj. length (km) 6.37 3.25 252.49
Max. traj. length (km) 80.61 99.41 1.16e5
Spatial area (km2) 16.0 $\times$ 20.1 9.9 $\times$ 9.6 1.0e3 $\times$ 1.2e3
[…]
Table 4: Dataset Statistics.
All experiments are run on a machine with an Intel Xeon Gold 6326 CPU, an NVIDIA A100 80 GB GPU and 64 GB RAM.
Appendix C Additional Experimental Results
This section includes a series of additional experimental results. While these results are not as central as those in the main text, they still provide solid evidence of the model’s effectiveness, robustness and efficiency. These include:
  1. C1.
    Impact of training set sizes
  2. C2.
    Evaluation on low quality data
  3. C3.
    Evaluation on test sets in unseen distributions
  4. C4.
    Additional results on ablation studies
    1. -
      Impact of self-attention modules
    2. -
      Embedding quality of the sub-view encoder
  5. C5.
    Additional results on parameter studies
    1. -
      Impact of batch sizes
    2. -
      Impact of $\lambda$
    3. -
      Impact of number of self-attention layers
  6. C6.
    Additional results of model training time
C.1 Impact of Training Set Size
We also investigate the impact of training set size on model accuracy. Each default dataset used in Section 5.2 contains 7,000, 1,000 and 2,000 trajectories for training, validation and testing, respectively. We increase the training set sizes to 14,000 and 28,000 while keeping the validation and test sets unchanged. These expanded datasets are referred to with “-2x” and “-4x” in Table 5, respectively.
In the table, we report the results of TSMini (repeating the experiments as before) comparing with the overall best baseline TrajCL to approximate Fréchet on Xian. TSMini consistently outperforms TrajCL, while additional training data does not lead to significant accuracy improvements in learning to approximate Fréchet. Notably, TSMini achieves higher accuracy on Xian compared to TrajCL does on Xian-4x, despite Xian-4x is a larger training set. Such results indicate that TSMini remains effective even with a smaller training set, making it highly practical.
[…]
Xian TrajCL 0.634 0.732 0.948
TSMini 0.812 0.890 0.997
Xian-2x TrajCL 0.726 0.811 0.987
TSMini 0.822 0.899 0.999
Xian-4x TrajCL 0.729 0.816 0.990
TSMini 0.824 0.900 0.999
[…]
Table 5: Impact of Training Set Size
C.2 Evaluation on Low Quality Data
We evaluate TSMini on particularly low-quality data to randomly mask and shift points in test trajectories, following the TrajCL paper Chang et al. [2023]. We randomly mask 40% and 80% of the points in test trajectories (denoted as $M_{40\%}$ and $M_{80\%}$ respectively) and/or apply random shifts of 100 or 500 meters to the points (denoted as $S_{100}$ and $S_{500}$ respectively). We report HR@10 for learning Fréchet on Xian in Table 6. TSMini consistently outperforms TrajCL by 25% on average. This is because TSMini models trajectories in continuous space, making it more generalizable than grid-based discrete methods.
[…]
TrajCL 0.634 0.632 0.583 0.545 0.470 0.502 0.452
TSMini 0.808 0.799 0.669 0.793 0.686 0.783 0.559
[…]
Table 6: Model effectiveness on low quality test trajectories in HR@10
C.3 Evaluation on Test Sets in Unseen Distributions
Following TrajCL Chang et al. [2023], we conduct experiments where models were trained on one dataset and directly tested on the other dataset. This is a challenging experimental setup, as the model is evaluated on a completely unseen trajectory distribution. It tests the model’s robustness and generalization ability.
We use Porto and Xian datasets. Table 7 show the results, where $A\rightarrow B$ denotes that the model is trained on dataset $A$ and tested on dataset $B$. TSMini consistently outperforms TrajCL by 46%. Besides, compared to the results of querying the original trajectories in Table 1, TSMini degrades by only 24% on average, while TrajCL drops by 36%, confirming the generalizability of TSMini.
[…]
DTW EDwP Fréchet DTW EDwP Fréchet
TrajCL 0.230 0.368 0.548 0.261 0.284 0.254
TSMini 0.509 0.621 0.633 0.494 0.619 0.687
[…]
Table 7: Model effectiveness on test sets in unseen distributions in HR@10
C.4 Additional Results on Ablation Study
Impact of Self-Attention Modules
We investigate the impact of different self-attention modules. TSMini adapts the self-attention module from Llama-2 Touvron et al. [2023]. We compare it with a model variant that uses the vanilla self-attention of Transformer Vaswani et al. [2017], denoted as TSMini-vanillaSA. We repeat the experiments of the ablation study and report the HR@10 results. Similar result patterns are recorded with the other metrics.
[…]
Porto TSMini 0.765 0.796 0.754
TSMini-vanillaSA 0.751 0.793 0.740
Xian TSMini 0.754 0.797 0.812
TSMini-vanillaSA 0.751 0.783 0.799
Germany TSMini 0.717 0.822 0.869
TSMini-vanillaSA 0.704 0.811 0.858
[…]
Table 8: HR@10 vs. the Self-attention Module in the Trajectory Encoder of TSMini.
Table C.4 shows the results. TSMini outperforms TSMini-vanillaSA consistently, showing that the adapted self-attention from Llama-2 is more effective than the vanilla self-attention for trajectory similarity learning. Meanwhile, the advantage is just 1.4% on average. This suggests that the main performance gain of TSMini compared with the existing models comes from the sub-view encoder and the $k$NN-guided loss, rather than the self-attention from Llama-2.
Embedding Quality of the Sub-View Encoder
[…]
TSMini 0.919 0.954 0.939
TSMini-w/o-S 0.903 0.935 0.929
[…]
Table 9: Embedding Quality of TSMini with and without Sub-View Encoder – Rand Index (RI) of Trajectory Clustering
Apart from the ablation study on SVEnc in Section 5.2, we further conduct experiments to investigate the effectiveness of SVEnc in capturing movement patterns.
We evaluate the clustering accuracy of TSMini (with sub-view-based inputs) and TSMiniw/o-S (using the cell-based inputs as in T3S and TrajCL) based on the generated trajectory embeddings, as this can indicate how SVEnc learns trajectory representations between similar trajectories. We follow the experimental settings used in a recent study Chang et al. [2024a], applying the $k$-medoids algorithm ($k$=10) and measuring cluster accuracy using the Rand Index (RI). RI (higher the better) measures the percentage of ground-truth similar trajectory pairs (based on non-learned measures) being correctly assigned to the same cluster. We report the results on Xian in Table 9.
We see that TSMini (using SVEnc) achieves higher RI consistently than the model variant using cell-based inputs, verifying that SVEnc helps TSMini produce more accurate trajectory representations.
C.5 Additional Results on Parameter Study
We report the full results of parameter study in this section. We investigate how (1) the batch size $N$, (2) the hyper-parameter $\lambda$ in the loss function, and (3) the number of self-attention layers #$\#layers$ in the trajectory encoder of TSMini impact model performance. We repeat the experiments as in Section 5.2 and only report HR@10. Similar patterns are observed on the other metrics.
(a) On Porto
(b) On Xian
(c) On Germany
Figure 7: HR@10 vs. the batch size ($N$ in Eq. (11)).
Impact of Batch Size
Figure C.4 shows the results on varying $N$. The HR@10 results generally increase as $N$ increases. This is expected, since a larger batch size (i.e., a larger $N$ in Eq. (11)) allows TSMini to learn more relative similarity values from the $k$NN-guided loss. We use a batch size of 128 by default to keep inline with the baselines for ease of comparison, although a larger batch size could further improve the accuracy of our model.
(a) On Porto
(b) On Xian
(c) On Germany
Figure 8: HR@10 vs. $\lambda$ in the loss ($\lambda$ in Eq. (13)).
Impact of $\lambda$.
Figure C.4 shows the results on varying $\lambda$ in Eq. (13). Except when $\lambda$ is 0 or 1, the value of $\lambda$ has a light impact on TSMini. Such results indicate that both loss terms contribute to the accuracy of TSMini, while using the $k$NN-guided loss alone ($\lambda=0$) is more effective than using the MSE loss ($\lambda=1$).
(a) On Porto
(b) On Xian
(c) On Germany
Figure 9: HR@10 vs. the number of self-attention layers
Impact of the Number of Self-attention Layers.
Figure C.4 shows the results on varying #$\#layers$. When #$\#layers$ increases, HR@10 first increases slightly and then decreases slightly. More layers initially help improve the model accuracy, as more learnable parameters are introduced, while an excessive number of parameters can lead to the overfitting problem and hence negatively impact the model accuracy. We use 1-layer TSMini by default for efficiency, as the model accuracy is already high with this setup.
C.6 Additional Results on Model Training Time
We further report the average training time per epoch and the maximum number of training epochs on the Xian dataset in Table 10, extending the results in Table 3.
It is worth mentioning that the reported total training time is end-to-end, including, e.g., preprocessing time. For example, TrajGAT preprocesses to convert raw trajectories into graphs, which has a high time cost. Besides, the maximum number of training epochs (i.e., Maximum #epochs) is not the actual number of epochs each model is trained for, as most models apply early stopping. Thus, the average training time per epoch has no direct correlation with the total time or the maximum number of epochs.
[…]
Total training
time
Average training
time per epoch
Maximum
#epochs
NeuTraj 0.12M 5,661s 283s 20
T3S 0.20M 2,885s 145s 20
TMN 0.18M 6,176s 601s 20
TrajGAT 2.84M 34,866s 1,212s 20
TrajCL 1.22M 3,006s 151s 30
KGTS 2.99M 5,600s 17s 300
TSMini 0.28M 2,605s 100s 40
[…]
Table 10: Additional Results on Model Training Time (Xian)